### v3

# Scalable Multi-Robot Motion Planning Using Guidance-Informed Hypergraphs

Abstract
In this work, we propose a method for multiple mobile robot motion planning that efficiently plans for robot teams up to an order of magnitude larger than existing state-of-the-art methods in congested settings with narrow passages in the environment. We achieve this improvement in scalability by adapting the state-of-the-art Decomposable State Space Hypergraph (DaSH) planning framework to expand the set of problems it can support to include those without a highly structured planning space and those with kinodynamic constraints. We accomplish this by exploiting guidance about a problem’s structure to limit exploration of the planning space and through modifying DaSH’s conflict resolution scheme. This guidance captures when coordination between robots is necessary, allowing us to decompose the intractably large multi-robot search space while limiting risk of inter-robot conflicts by composing relevant robot groups together while planning.
I Introduction
Multi-robot systems are becoming increasingly prevalent in settings like warehouses and factories. These often constrained environments require careful yet computationally fast motion planning for each robot to accomplish its task without collision. Some navigation approaches compute individual decoupled paths for each robot and resolve conflicts online [1]; however, in congested environments, this introduces the risk of deadlock. Alternatively, coupled offline multi-robot motion planning (MRMP) approaches [1] plan for all robots together in the composite space, the joint planning space of all robots. While these methods provide the high level of coordination required to solve difficult problems, due to the large size of the search space, they are computationally intractable for large robot teams. Hybrid methods [2, 3, 4] seek to combine the benefits of decoupled and coupled planning, providing an increased level of coordination when necessary while also maintaining scalability.
Figure 1: A comparison of the DaSH framework and the Guided-DaSH framework. In DaSH, the task space hypergraph and motion hypergraph are exhaustively constructed. The motion hypergraph is then queried to find an optimistic schedule which then undergoes conflict resolution. In Guided-DaSH, a heuristic search over the implicit problem space allows us to build a minimal portion of the task space hypergraph and motion hypergraph. Because of this, we can directly extract the motion solution from the motion hypergraph.
The Decomposable State Space Hypergraph (DaSH) [5] framework presents a general structure for hybrid planning that benefits from a sparse representation of the planning space. This representation models when coordination between robots is needed, allowing the planner to stay in low-dimensional search spaces when possible, and captures the connections between these spaces. The framework features a hierarchical approach consisting of a high-level representation of the planning space, the task space hypergraph, that directs construction of a motion hypergraph from which an optimistic motion solution which is not guaranteed to be collision-free is found and then validated. The DaSH framework can be applied to develop methods for a variety of multi-robot planning problems. Here, we propose leveraging DaSH for problems including mobile robot motion planning in environments with narrow passages, a classical challenge in motion planning exacerbated in the multi-robot setting, where the robots may also have kinodynamic constraints which further complicate planning.
DaSH proposes an approach to exhaustively constructing the task space hypergraph that performs well for problems like multi-manipulator rearrangement [5], where the search space has an exploitable structure since actions can be classified as pick, place, handoff, etc., but becomes intractable for problems which lack this highly structured search space, such as mobile robot motion planning. Our insight is that we can leverage information about the planning problem to induce this structure while building the high-level representation and further exploit this guidance, in conjunction with a modified conflict resolution scheme, to efficiently search the planning space and support kinodyanmic constraints. In the case of mobile robot motion planning, we can effectively use workspace information to provide this guidance [6, 7, 8, 9, 10]. As such, we propose Guided-DaSH, our adaptation to the DaSH framework to support a wider range of problems and Workspace Guided-DaSH, an application of our framework to the problem of multiple mobile robot motion planning in environments with narrow passages. Our contributions include:
  • An adaptation of the DaSH framework to multi-robot planning problems without a highly structured search space and a novel hybrid MRMP method built on our adapted framework that leverages the workspace to guide planning in multiple ways.
  • Support for a broader range of problems, including planning under kinodynamic constraints within our framework via an adapted conflict resolution structure.
  • An extensive experimental validation showing that our novel approach scales to robot groups up to 128, an order of magnitude larger than previous methods.
II Preliminaries and Prior Work
In this section, we describe relevant motion planning preliminaries and discuss prior work in the field of multi-robot motion planning.
II-A Motion Planning Preliminaries
In motion planning, a configuration refers to a set of values for a robot’s degrees of freedom, including their position in the workspace, orientation, and other parameters such as joint angles. The configuration space ($\mathcal{C}_{space}$) consists of all possible configurations and can be partitioned into $\mathcal{C}_{free}$ and $\mathcal{C}_{obst}$ made up of valid and invalid configurations respectively. Motion planning is the problem of finding a path from a start configuration $q_{s}$ to a goal configuration $q_{g}$ through $\mathcal{C}_{free}$. Computing the full $\mathcal{C}_{obst}$ is intractable [11, 12], thus state-of-the-art methods turned to sampling-based motion planning [13, 14]. Two commonly used methods, Rapidly-exploring Random Trees (RRT) [14] and Probabilistic Roadmaps [13] build tree-based and graph-based representations of $\mathcal{C}_{free}$ respectively.
These methods have also been extended to problems which require satisfying kinematic constraints that limit the robot’s position, and dynamic constraints that limit the velocity or acceleration of the robot. Kinodynamic RRT [15] extends RRT to work with kinodynamic constraints through extensions of randomly sampled control inputs for random durations until a goal region is reached.
II-B Multi-robot Motion Planning
The MRMP problem is an extension of the motion planning problem which consists of finding valid paths $\Pi=\{\pi_{r_{1}},...,\pi_{r_{n}}\}$ for a set of robots $R=\{r_{1},...,r_{n}\}$ that also avoid inter-robot collision. Multi-agent pathfinding (MAPF) is the discrete state space equivalent of the MRMP problem which consists of finding collision-free paths for a set of robots over a given graph representation (e.g. a grid) rather than through continuous space. In environments with narrow passages, it is non-trivial to construct a graph representation that captures the connectivity of the planning space, hence our consideration of MRMP.
Some coupled methods, including Composite PRM [1] and Composite RRT [14] provide a high level of coordination by planning directly in the composite space, $C_{0}\times...\times C_{n-1}$ where $C_{i}$ is the configuration space of robot $i$. This becomes computationally intractable due to the high dimensionality of the space with large groups of robots. Other coupled methods, including MRdRRT [16] and its variants [17, 18], construct single-robot roadmaps and then use these to search an implicit composite space roadmap.
Hybrid methods combine decoupled and coupled planning. CBS-MP [2] performs a low-level search to find decoupled paths for each robot and resolves conflicts using a high-level search in the composite space. Several hybrid MAPF methods have been proposed to find collision-free paths given a representation of the $\mathcal{C}_{space}$. M* [3] conducts an A*-like search using a technique called subdimensional expansion where optimal individual paths for each robot are followed until a conflict occurs. Then, the conflicting agents are merged together and a new conflict-free composite path is found. Subdimensional Expansion RRT (sRRT) [19] applies subdimensional expansion to MRMP. It constructs individual policies by growing RRTs in the $C_{space}$ of each robot backward from the goal. It then expands a tree forward in the composite space by following the individual policies until conflict occurs. Then, new configurations are sampled for expansion in the composite space of the conflicting robots. While these methods perform well in open environments, in constrained settings, they expend excess computation on conflict resolution, limiting planning efficiency. Adaptive Robot Coordination (ARC) [20] first plans decoupled paths for each robot and then resolves conflicts by solving local subproblems in the composite spaces of the affected robots. Though it does focus computation within constrained settings, it does so reactively and spends excess time adjusting the local problem size to find the right resolution.
Kinodynamic Conflict-Based Search (K-CBS) [21] extends hybrid MRMP algorithms to accommodate kinodynamic constraints, drawing on the concepts of CBS-MP [2]. This approach introduces kinodynamic conflicts and employs a kinodynamic low-level planner akin to the kinodynamic RRT in [15], enabling the ability to solve high-dimensional problems with general dynamics.
II-C Guided Motion Planning
Guided motion planning methods leverage external information to efficiently find paths. Prior work has explored using knowledge of the workspace in the form of a wavefront expansion [9] or workspace decomposition regions [10] to guide single-robot planning. Many approaches use workspace skeletons, embedded graphs in which edges represent free areas of the workspace and vertices represent connections between them. Examples include medial axis skeletons [22] for 2D environments and mean curvature skeletons [23] for 3D. They are generally quick to compute, for example, medial axis skeletons can be computed in $O(n\log n)$ time where $n$ is the number of obstacle edges [24]. Skeleton-guided motion planning methods [7, 8] use dynamic sampling regions, bounded local areas of the workspace that advance along skeleton edges, to efficiently explore free areas of the environment, including narrow passages, until a path is found.
Prior work extends skeleton guidance to MRMP for scenarios with narrow passages which require a high level of coordination during planning. Composite Dynamic Region-biased RRT (CDR-RRT) [6] is a coupled method that lazily builds and searches over a composite skeleton, which is the Cartesian product graph of the skeleton for each robot in the group. MAPF is used to generate a path over the skeleton for each robot, taking into account each edge’s capacity as given by the clearance to obstacles in the environment. These paths are combined to form a path through the composite skeleton. Dynamic sampling regions advance over the composite skeleton edges, growing a tree until the goal is reached. Although CDR-RRT shows a significant improvement in scalability compared to other state-of-the-art methods in environments with narrow passages where coordination between the full robot group is required [6], considering the full composite space of all robots incurs significant computational overhead. We propose a hybrid method extended from CDR-RRT that uses topological guidance to inform when planning in the composite space (of all or a subgroup of robots) is necessary.
III Guided-DaSH
In this section, we describe our adaptation to the DaSH framework, Guided-DaSH, and explain the modifications we make to support motion planning problems without a highly structured search space and kinodynamic constraints. For a detailed description of the DaSH framework, see [5]. Figure 1 shows a comparison of Guided-DaSH and the original DaSH framework which we will elaborate on below. Algorithm 1 shows an overview of our framework.
Algorithm 1 Guided-DaSH
1:Heuristic $H$, Query $Q$
2:$\Pi\leftarrow\emptyset$triangleright Robot motion paths
3:$C\leftarrow\emptyset$triangleright Heuristic search constraints
4:while $\Pi=\emptyset$
5:    $\mathcal{H}_{ts}\leftarrow\textsc{ConstructTaskSpaceHypergraph}(H,C,Q)$
6:    $\mathcal{X}_{ts}\leftarrow\textsc{OrderByDependency}(\mathcal{H}_{ts})$
7:    $\mathcal{H}_{m}\leftarrow\emptyset$triangleright Initialize motion hypergraph
8:    for $x_{ts}\in\mathcal{X}_{ts}$
9:         $\pi\leftarrow\textsc{ComputePath}(G,x_{ts})$
10:         if $\pi=\emptyset$
11:             $C\leftarrow C\cup\textsc{GetConstraint}(x_{ts})$
12:             break $\triangleright$ Restart heuristic search
13:         $\mathcal{X}_{conf}\leftarrow\textsc{FindCollisions}(\mathcal{H}_{m},\pi)$
14:         if $\mathcal{X}_{conf}\neq\emptyset$
15:             $C\leftarrow C\cup\textsc{GetConstraints}(\mathcal{X}_{conf}\cup\{x_{ts}\})$
16:             break $\triangleright$ Restart heuristic search
17:         $\mathcal{H}_{m}.\textsc{AddPath}(\pi)$
18:    $\Pi\leftarrow\textsc{ExtractMotionPaths}(\mathcal{H}_{m})$
III-A Task Space Hypergraph Construction
In the task space hypergraph $\mathcal{H}_{ts}=\{\mathcal{V}_{ts},\mathcal{E}_{ts}\}$, vertices $\mathcal{V}_{ts}$ are task space elements which represent local planning spaces and are composed of the robots required to coordinate at that point and relevant constraints they should adhere to. Hyperarcs $\mathcal{E}_{ts}$ can be categorized as composition, representing changes in the robots considered and constraints, or transition, representing changes in constraints for a fixed set of robots. The original DaSH framework [5] exhaustively constructs $\mathcal{H}_{ts}$; however, by using guidance to inform which task space elements are needed, we can efficiently construct a minimal portion of $\mathcal{H}_{ts}$. This makes our adapted framework advantageous for problems where exhaustive construction is intractable.
To construct a minimal portion of the task space hypergraph, we require guidance in the form of domain knowledge about the problem being considered. This defines the set of constraints which can exist in a task space element. Additionally, this knowledge must inform a heuristic $H$ required to search over the implicit problem space and construct a series of task space elements $\mathcal{V}_{ts}$ that satisfy a given query. Hyperarcs $\mathcal{E}_{ts}$ are added between subsequent task space elements to form $\mathcal{H}_{ts}$. The heuristic must also take as input a set of constraints that represent prohibited task space elements that cannot exist in the solution since they have failed to be translated into a motion solution. Searching over the implicit problem space can be achieved by searching over a smaller representative space that captures when coordination between robots is required. We provide an example in which information about the workspace topology is used to guide composition decisions in Section IV.
III-B Motion Hypergraph Construction
The motion hypergraph $\mathcal{H}_{m}$ stores the configurations and motions between them that make up the robots’ motion paths through the task space elements and hyperarcs from $\mathcal{H}_{ts}$. It is often necessary to leverage a guided planning method to ensure that the constraints associated with each task space element are met when constructing local paths. We detail an example of such a method in Section IV.
For each task space element and transition hyperarc in $x_{ts}\in\mathcal{H}_{ts}$, ordered from start to goal, a local path is computed in the composite space of the robots contained therein. These paths become vertices and motion hyperarcs in $\mathcal{H}_{m}$. Transitions often require a different method of path construction to account for, e.g., grasp constraints or other problem-specific factors. It is possible that constructing a local path for a task space element or transition could fail. In this case of a task space element failing, a constraint is added to the heuristic search to prohibit that task space element from existing in $\mathcal{H}_{ts}$. Similarly, in the case of a transition failing, constraints are added to prohibit the task space elements in the corresponding incoming and outgoing sets of the transition from existing simultaneously in $\mathcal{H}_{ts}$.
In the original DaSH formulation, the entire motion hypergraph is constructed before a scheduled query is performed that finds and resolves conflicts between robot paths through different task space elements or transitions. In our adapted framework; however, since we know from the heuristic search which task space element’s motions should occur concurrently, we can interleave motion hypergraph construction and conflict detection. By finding conflicts as they occur rather than after the motion hypergraph is constructed, we minimize wasted computation and can directly extract the motion solution from the motion hypergraph, bypassing the query stage.
Our modification to the conflict resolution stage also allows us to support kinodynamic planning. The scheduled query proposed in the original DaSH framework resolves conflicts by inducing waiting to prevent robots from colliding with each other. When planning under kinodynamic constraints, we cannot assume the ability of robots to instantaneously change velocity. Additionally, to avoid solving computationally difficult optimal control point-to-point problems, we cannot reuse local paths for task space elements occurring after conflicts in the robots’ paths. Thus, we must replan affected paths to avoid conflicts.
As we construct $\mathcal{H}_{m}$, after each local path is computed, we check for collisions with existing local paths occurring simultaneously by considering intermediate configurations along the paths. When a conflict is found, constraints are added to the heuristic search to prevent the corresponding task space elements or transitions from existing in $\mathcal{H}_{ts}$ together. When constraints are added, the heuristic search is redone, $\mathcal{H}_{ts}$ is recomputed, and construction of $\mathcal{H}_{m}$ restarts. This process continues until a conflict-free motion solution is found.
IV Workspace Guided-DaSH
In this section, we explain how we apply the Guided-DaSH framework for mobile robot MRMP in congested environments with narrow passages to propose the Workspace Guided-DaSH method and describe how we leverage the workspace skeleton to guide both the construction of $\mathcal{H}_{ts}$ and its translation into $\mathcal{H}_{m}$.
IV-A Task Space Hypergraph Construction
Considering mobile robots, coordination between the motions of robots is required when they are physically in proximity to each other. Our insight is that we can leverage the skeleton to approximate when coordination is necessary by modeling the movement of the robots through the workspace as movement along the skeleton. Specifically, coordination is necessary when robots are traversing the same skeleton edge (either in the same or opposite direction) or traveling through the same skeleton vertex. This formulation imposes a structure on the search space that we capture through our heuristic search.
Workspace Guided-DaSH follows the structure shown in Alg. 1. Given a skeleton of the workspace, we first conduct our heuristic search to find a path over the skeleton for each robot using a MAPF algorithm. We set the capacity of each skeleton vertex and edge as given by the minimum clearance to an obstacle. We then combine these paths to form the minimal necessary portion of the task space hypergraph. Fig. 2 shows an example of the conversion of the MAPF solution into $\mathcal{H}_{ts}$. Within task space elements, constraints represent the skeleton segments that each robot should traverse at that stage. Robots traversing the same skeleton edge are grouped into the same task space element using composition hyperarcs. As robots move away from each other onto different edges, they are decoupled into separate task space elements, boosting planning efficiency by considering smaller $\mathcal{C}_{spaces}$. Transition hyperarcs capture robots moving through the same skeleton vertex as well as robots moving between segments of the skeleton (i.e. moving between task space elements with the same robots but different constraints).
(a) Segment of Robot Paths over Skeleton
(b) Generated Portion of Task Space Hypergraph
Figure 2: A segment of the robot paths over the skeleton (gray) and the corresponding portion of the task space hypergraph. Each task space element shows the corresponding set of robots and the relevant portion of the topological skeleton. In (b), arrows indicate robots moving along a skeleton edge and dots indicate a decoupled state. Solid lines indicate transition hyperarcs and dashed lines represent composition hyperarcs. The green and red robots are initially moving in opposite directions along the same skeleton edge. Then, red and blue reach the skeleton vertex together and go on to different edges. The yellow robot begins moving in the opposite direction of the red robot when the red robot reaches the next skeleton edge.
IV-B Translation to Low-level Motion Hypergraph
Once $\mathcal{H}_{ts}$ is constructed, we compute a set of paths to form $\mathcal{H}_{m}$. To ensure that the robots roughly follow their paths over the skeleton and that narrow passages in the environment are efficiently traversed, we leverage the skeleton to guide the construction of local paths along skeleton edges. To compute paths through task space elements, we use the dynamic sampling region advancement procedure from CDR-RRT [6].
Considering transition hyperarcs that represent multiple robots moving through a skeleton vertex, we stop the movement of the robots along the incoming skeleton edges a small distance $\delta$ from the end of each edge. We’ve found that setting $\delta$ to the diameter of the region used for the dynamic sampling region advancement works well. We then grow an RRT [14] in the composite space of these robots from the last configuration on the robots’ incoming skeleton edges to a sampled configuration at least $\delta$ distance away from the skeleton vertex of the robots’ next skeleton edges. Here, we simply grow an RRT rather than using CDR-RRT to avoid forcing the robots to reach the skeleton vertex at the same time, creating congestion. Composition hyperarcs and transition hyperarcs not associated with movement in the task space hypergraph correspond directly to vertices and hyperarcs in the motion hypergraph.
We consider construction of a local path to have failed if the RRT fails to be extended a given number of times during planning. In this case, we impose a constraint that the MAPF solution from which we build $\mathcal{H}_{ts}$ cannot contain the same group of robots traversing along either a skeleton edge that failed or the same groups of robots traversing along the incoming and outgoing skeleton edges to a skeleton vertex that failed. We then recompute $\mathcal{H}_{ts}$ and try again to construct $\mathcal{H}_{m}$. Similarly, if conflicts are found between robots traversing different portions of the skeleton simultaneously, constraints are imposed to prevent the relevant robot groups from traversing these segments of the skeleton together.
IV-C Kinodynamic Planning
Incorporating kinodynamic constraints into Workspace Guided-DaSH is straightforward due to the modularity of the framework. As kinodynamic constraints influence the motion of the robots, it is imperative that the low-level motion hypergraph, $\mathcal{H}_{m}$, incorporates the kinodynamic constraints during construction. To this end, we adapt the use of CDR-RRT, employed during the construction of $\mathcal{H}_{m}$ to incorporate kinodynamic constraints. This is achieved by substituting their underlying RRT with Kinodynamic RRT [15].
A control input is generated by sampling the control space uniformly at random which is then integrated forward during the RRT extension phase for a randomly sampled duration. In order to utilize the guidance provided by the CDR-RRT, we generate and integrate forward many random control inputs and select the input which moves the system closest to the dynamic sampling region. The kinodynamic RRT continues to expand until either a maximum number of states have been added or the goal region has been reached. We refer the readers to [15] for more details on kinodynamic RRT.
(a) Warehouse
(b) Tunnels Environment
(c) Mining
(d) GridMaze
(e) GridMaze Query
Figure 3: (a) In the Warehouse environment, robots located on the top and bottom of each narrow aisle must swap places with the robot they are vertically aligned with. (b) In the Tunnels environment, robots must find and traverse multiple tunnels while moving from start positions in the back tunnel to goal positions in the front tunnel. (c) In the mining environment, robots must swap places with another robot located in an adjacent mine shaft, robot start positions are located at the front of the tunnels and at intersections between tunnels. (d-e) In the GridMaze environment, robots must move between randomly located starts (red) and goals (blue) within the maze, some of which overlap.
V Validation
We run scaling MRMP scenarios in environments designed to highlight the strengths and weaknesses of Workspace Guided-DaSH (abbreviated WG-DaSH in results) relative to other state-of-the-art approaches.
V-A Experimental Setup
We measure the performance of Workspace Guided-DaSH against CDR-RRT [6] as well as a set of other approaches that are relevant for comparison. Composite RRT [14] operates in the full composite space, providing a high level of coordination. MRdRRT [16] was originally intended for manipulators, but its implicit search of the composite space makes it pertinent for comparison. sRRT [19] uses subdimensional expansion to limit the size of the search space which is similar in concept to our hypergraph representation. ARC [20] composes robots into higher dimensional planning spaces based on encountered conflicts, whereas our method uses workspace guidance to preemptively avoid likely conflicts. Thus, comparing our method against ARC effectively measures the impact of this guidance. We use Conflict-based Search [25] over the workspace skeleton to perform the heuristic search for Workspace Guided-DaSH. Each method was implemented in the open-source Parasol Planning Library. Holonomic experiments were run on a 4-core CPU (4GHz, 32GB RAM), while kinodynamic experiments used a 24-core CPU (6GHz, 64GB RAM). We run 15 seeds for each scenario. Each method is given 600 seconds to solve a scenario or is considered a failure. We report planning time as well as path cost given by the makespan.
TABLE I: Warehouse results. Data omitted where all seeds failed
Method Robots Runtime (s) Cost (s) Success
Avg Std Avg Std
WG-DaSH 2 0.7 0.3 31.6 4.4 100%
CDR-RRT 1.5 0.8 22.3 2.6 100%
ARC 9.6 15.3 22.4 5.3 100%
sRRT 3.4 3.7 29.8 9.5 100%
Comp. RRT 22.6 24.2 25.7 12.2 100%
MRdRRT 21.6 26.4 82.1 35.2 53.3%
  WG-DaSH 4 1.8 0.6 32.8 3.2 100%
CDR-RRT 2.4 1.0 27.6 6.7 100%
ARC 16.7 18.9 22.8 7.9 93.3%
sRRT 14.7 9.6 33.2 18.9 86.7%
Comp. RRT 274.0 - 19.2 - 6.7%
MRdRRT 295.0 233.4 107.1 77.4 20.0%
  WG-DaSH 8 4.6 3.2 32.5 3.0 100%
CDR-RRT 23.0 8.2 42.2 7.5 100%
ARC 37.7 49.6 21.0 0.9 80.0%
sRRT 61.4 40.1 20.9 0.3 13.3%
  WG-DaSH 16 9.5 2.1 31.4 3.4 100%
CDR-RRT 288.1 100.2 50.1 7.1 93.3%
ARC 97.1 188.5 20.7 0.6 60.0%
  WG-DaSH 32 29.5 5.8 34.3 4.0 100%
ARC 72.4 7.1 20.9 0.7 26.6%
  WG-DaSH 64 54.1 11.4 35.6 2.7 100%
  WG-DaSH 128 250.7 29.4 36.6 5.0 60.0%
V-B Environments
Here, we describe our experimental environments and explain which aspects of our approach these scenarios are designed to highlight.
V-B1 Warehouse (Fig. 3(a))
The Warehouse environment features several narrow aisles through which robots must swap places with the robot opposite them. In the holonomic scenario, the width of the narrow passages is approximately 2.5 times the diameter of the disk robots we consider. We additionally run a kinodynamic variant with doubled aisle widths. This scenario measures each approach’s ability to efficiently handle narrow passages within the environment. We highlight Workspace Guided-DaSH’s ability to decompose the problem using topological guidance.
V-B2 Tunnels (Fig. 3(b))
The Tunnels environment again features several narrow passages. Here, because of the 3D workspace, the planning space becomes very large. This scenario is intended to measure each algorithm’s performance with respect to the geometric complexity, rather than the congestion, of the environment. We consider L-shaped robots with the width of their longest dimension slightly less than half the width of the narrow passages. Each robot must find and traverse multiple narrow passages to reach its goal, but its shortest path does not conflict with any other robot’s.
TABLE II: Tunnels results. Data omitted where all seeds failed
Method Robots Runtime (s) Cost (s) Success
Avg Std Avg Std
WG-DaSH 2 4.2 0.9 166.0 5.1 100%
CDR-RRT 6.2 1.1 140.6 11.2 100%
sRRT 129.6 175.0 175.3 26.1 40.0%
  WG-DaSH 4 8.2 1.0 158.3 9.0 100%
CDR-RRT 17.3 3.3 152.9 16.4 100%
sRRT 327.0 - 161.7 - 6.7%
  WG-DaSH 8 14.1 2.4 159.8 8.4 100%
CDR-RRT 36.0 6.8 156.1 9.0 100%
  WG-DaSH 16 22.5 1.9 164.8 9.4 100%
CDR-RRT 73.1 8.5 163.9 16.7 100%
  WG-DaSH 32 54.0 6.2 163.0 8.0 100%
CDR-RRT 252.6 10.0 159.8 4.7 93.3%
  WG-DaSH 64 132.4 7.0 163.1 6.0 100%
V-B3 Mining (Fig. 3(c))
The Mining environment is a 3D environment meant to mimic a set of mine shafts. Again, it features several narrow passages; however, here robots must move in close proximity to each other. We use the same robots as the Tunnels environment. We show that Workspace Guided-DaSH is capable of efficiently finding paths for large robot teams in problems representative of real-world scenarios.
V-B4 GridMaze (Fig. 3(d))
The GridMaze environment is a 3D environment featuring several intersecting narrow tunnels intended to elicit high congestion for large groups of robots in addition to geometric complexity. We use a rectangular prism robot whose smallest dimension is slightly less than half the width of the narrow passages. We demonstrate Workspace Guided-DaSH’s ability to efficiently find paths for large groups of robots in these congested settings.
TABLE III: Mining results. Data omitted where all seeds failed
Method Robots Runtime (s) Cost (s) Success
Avg Std Avg Std
WG-DaSH 2 1.3 0.3 125.6 6.5 100%
CDR-RRT 1.9 0.4 113.0 11.7 100%
sRRT 146.8 161.7 131.5 17.5 33.3%
  WG-DaSH 4 2.4 0.5 122.6 10.6 100%
CDR-RRT 16.0 6.1 151.2 19.5 100%
  WG-DaSH 8 5.0 0.8 132.2 6.7 100%
CDR-RRT 45.7 14.8 160.6 12.9 100%
  WG-DaSH 16 11.0 1.0 130.3 5.8 100%
CDR-RRT 118.9 44.3 179.0 17.5 100%
  WG-DaSH 32 28.1 3.5 127.6 6.7 100%
  WG-DaSH 64 80.2 6.9 124.5 4.5 86.7%
V-C Experimental Results
In the Warehouse scenario (results in Tab. I), all methods except MRdRRT achieve a 100% success rate in the 2-robot scenario. However, the methods that do not leverage topological guidance begin to struggle with the 4-robot scenario and multiple cannot solve the 8-robot scenario. Of those, only ARC is able to solve up to the 32-robot scenario; however, with a success rate of only 26.6%. CDR-RRT and sRRT are able to solve up to the 8-robot scenario where the exponential size of the composite space leads to a significant decrease in performance. Workspace Guided-DaSH, like CDR-RRT, sees the benefit of using the skeleton representation to guide planning through narrow passages and the coordination provided by using the composite space of multiple robots. Additionally, by using the workspace to inform the planner when to move between different levels of composition, Workspace Guided-DaSH is able to focus planning within smaller spaces than CDR-RRT, making it able to plan for significantly larger robot teams, up to 128 robots, within the 600-second time limit.
In the Tunnels environment (results in Tab. II), the size of the search space is increased significantly relative to the Warehouse environment since the robot now has six degrees of freedom. Composite RRT, MRdRRT, and ARC are unable to find a solution for the 2-robot scenario due both to the large size of the search space and the complexity of the environment. sRRT performs better because of its ability to decompose the search space, but struggles with the narrow passages and only one seed was able to solve the 4-robot scenario. Here, the narrow passage problem is exacerbated since the query requires the planner to find and traverse multiple narrow passages. By leveraging skeleton guidance, CDR-RRT and Workspace Guided-DaSH are able to find solutions for larger groups of robots. Searching the composite space hinders CDR-RRT’s performance, leaving it unable to solve the 64-robot scenario within the time limit. Workspace Guided-DaSH is able to efficiently find solutions for up to 64 robots within the time limit due to its ability to decompose the planning space.
TABLE IV: GridMaze results. Data omitted where all seeds failed
Method Robots Runtime (s) Cost (s) Success
Avg Std Avg Std
WG-DaSH 2 0.3 0.04 44.0 5.5 100%
CDR-RRT 1.0 0.3 63.9 8.3 100%
ARC 146.4 78.2 25.7 2.0 100%
sRRT 4.7 4.8 30.7 3.8 100%
Comp. RRT 425.0 136.0 59.4 11.5 20.0%
  WG-DaSH 4 1.0 0.5 44.2 3.8 100%
CDR-RRT 46.0 17.7 120.6 15.9 100%
ARC 215.2 137.6 26.3 2.1 93.3%
sRRT 19.2 6.0 30.9 4.1 100%
  WG-DaSH 8 3.2 1.5 42.3 3.3 100%
ARC 361.4 144.2 26.4 2.6 93.3%
sRRT 25.5 10.6 32.8 9.7 100%
  WG-DaSH 16 3.8 1.6 44.2 4.1 100%
ARC 437.0 55.9 26.0 1.9 40.0%
sRRT 38.1 10.8 29.3 2.6 100%
  WG-DaSH 32 25.5 7.9 36.3 16.4 93.3%
In the Mining environment (results in Tab. III), Composite RRT and MRdRRT are unable to find a solution for the 2-robot scenario due to the large size of the search space and inter-robot conflicts. Similar to the Tunnels scenario, ARC’s performance is impacted by the environment complexity since it does not use topological guidance and it fails to find solutions for the 2-robot scenario within the time limit. sRRT struggles to find solutions for the 2-robot scenario and cannot solve the 4-robot scenario. By leveraging topological guidance to limit exploration of the composite space, CDR-RRT is able to solve up to the 16-robot scenario. However, it is unable to solve the 32-robot scenario. Workspace Guided-DaSH, by considering the composite space only when necessary, is able to efficiently solve up to the 64-robot scenario.
In the 3D GridMaze scenario (results in Tab. IV), as a result of the large size of the planning space as well as the difficulty posed by the long curved narrow passages, Composite RRT struggles with the 2-robot scenario and fails to produce a solution for the 4-robot scenario. MRdRRT fails to solve the 2-robot scenario. While CDR-RRT is able to solve the 4-robot scenario, the large size of the composite space prevents it from finding a solution for the 8-robot scenario within the time limit. sRRT and ARC benefit from searching a lower-dimensional space when robots are not in conflict, but fail to solve the 32-robot scenario where conflicts become prevalent due to the high level of congestion of the environment. Similar to sRRT, Workspace Guided-DaSH searches lower-dimensional spaces, however, our method also uses the workspace skeleton for conflict resolution. When a conflict is too difficult to be resolved quickly via random sampling, we impose constraints on the robots’ abstract paths through the workspace which allows us to find conflict-free paths more efficiently. Thus, Workspace Guided-DaSH is able to find solutions up to 32 robots. Although this replanning mechanism sometimes results in slightly higher solution costs relative to other methods, a post-processing trajectory optimization step is generally used in practice to smooth paths.
V-D Kinodynamic Experiments
To evaluate Workspace Guided-DaSH with kinodynamic constraints, we use Kinodynamic Workspace Guided-DaSH (K-WG-DaSH) incorporating the modifications in Section IV-C. Kinodynamic-CBS [21] (K-CBS) is utilized as our baseline kinodynamic planner as it uses the same Kinodynamic RRT we employ in K-WG-DaSH, however they utilize a merge bound [26] creating composite robots when more coordination is needed. This will allow us to evaluate the benefits of the coordination guidance provided by K-WG-DaSH. Each method was given $45$ minutes to find a solution to planning problems in the warehouse environment, described in Section V-B1, using robots with $2$nd-order car dynamics.
In the kinodynamic warehouse environment (results in Tab.  V), K-WG-DaSH finds solutions for problems with up to 6 robots consistently while K-CBS begins to struggle at 4 robots. In the 2 robot scenario, K-WG-DaSH maintains a high success rate with a lower planning time at the expense of a higher path cost as opposed to K-CBS which is inconsistent at finding solutions with a higher planning time but with a lower path cost. We attribute the scalability to the topological and coordination guidance provided by K-WG-DaSH.
TABLE V: Kinodynamic Warehouse results. Data omitted where all seeds failed
Method Robots Runtime (s) Cost (s) Success
Avg Std Avg Std
K-WG-DaSH 2 201.2 119.1 46.7 9.9 93.3%
K-CBS 390.1 365.9 34.4 2.3 46.7%
  K-WG-DaSH 4 580.2 376.2 41.9 7.3 93.3%
K-CBS 445.0 473.3 35.1 2.7 20.0%
  K-WG-DaSH 6 1261.6 524.5 32.1 14.8 86.7%
K-CBS 56.8 - 33.5 - 6.7%
  K-WG-DaSH 8 2147.2 313.1 19.1 16.0 33.3%
  K-WG-DaSH 10 2544.7 - 36.5 - 6.7%
VI Conclusion and Future Work
In this work, we present both the Guided-DaSH framework for hybrid multi-robot planning and Workspace Guided-DaSH, an application of this framework to mobile robot MRMP in congested environments. Our adapted framework builds off of a state-of-the-art hybrid planning framework, DaSH, using guidance to enable its use for a wider range of multi-robot planning problems which lack a highly structured search space. Additionally, we add support for kinodynamic planning via a modified conflict resolution structure, which also boosts the scalability of our framework by eliminating unnecessary work when finding motion solutions.
Workspace Guided-DaSH leverages topological guidance in multiple ways, both to move between different levels of composition during planning and to guide sampling through narrow passages in the workspace. By using knowledge of the workspace, our method is able to efficiently find paths for robot groups of size up to an order of magnitude larger than existing state-of-the-art methods. Building off of this work, we plan to explore the use of different kinodynamic low-level motion planners which plan more intelligently for multi-robot systems by interleaving sampling-based planning and trajectory optimization [27]. Additionally, we plan to explore using parallel computing to further improve the performance of this method, taking advantage of the hypergraph’s ability to naturally decompose the multi-robot planning problems.